# Exploring New Physics with PandaX-4T Low Energy Electronic Recoil Data Xinning Zeng

### v1

@@
zhou˙xp@buaa.edu.cnSchool of Physics, Beihang University, Beijing 102206, China
@@ Abstract
We use the large liquid xenon detector PandaX-4T to search for novel electronic recoil signals induced by solar axions, neutrinos with anomalous magnetic moment, axion-like particles, dark photons, and light fermionic dark matter.
A detailed background model is established with the latest datasets with 1.54 exposure.
@@
In addition to the traditional DM-induced nuclear recoil (NR) events, particles from many new physics models can also produce novel electronic recoil signals (NERS), which have attracted increasing attention recently [6, 7, 8, 9].
Axions [10, 11, 12] can be produced in the Sun via the so-called ABC effect (Atomic recombination and de-excitation, Bremsstrahlung, and Compton scattering processes) [13, 14], the Primakoff conversion [15], and the nuclear magnetic transitions [16], related to the axion-electron coupling , axion-photon coupling , and effective axion-nucleon coupling , respectively.
The Sun also produces a vast number of neutrinos, which may carry anomalous magnetic moment due to the Majorana nature of neutrino or exotic new physics beyond the Standard Model [17].
With their kinetic energy governed by the temperature of the Sun, these particles can produce electronic recoil (ER) events with a typical energy from keV to a couple of tens of keV.
Within this energy range, alternative light dark matter models such as Axion-Like Particles (ALPs), dark photons [18], or light fermionic DM converting into active neutrinos [19] can also manifest themselves with characteristic ER spectra.
This paper reports an updated NERS search using the low energy ER data below 30 keV from PandaX-4T, a LXe detector that operates in the China Jinping Underground Laboratory (CJPL),
with a total live exposure of 1.54 tonneyear.
@@
For brevity, we leave detailed production and detection cross-sections and expected NERS spectra in the Supplementary Materials.
To set the stage, the signal of ABC solar axions and a 170 keV/ light fermionic DM absorption are shown in Fig. 1, with featured spectra located at the low and high end of the region-of-interest (ROI).
The keys to this analysis lie in the understanding of the detector responses, as well as robust prior knowledge of the background using data outside the ROI region.
Figure 1: Theoretical (shaded) and expected energy spectra (solid for Run0, dashed for Run1) for ABC axion with a coupling =3.1 (red shaded for theoretical, yellow unshaded for expected) and light fermionic DM- conversion with = 170 keV and cm2 (blue shaded for theoretical, purple unshaded for expected).
@@
Table 1: Operating conditions of Run0 and Run1. Duration represents the calendar days and live time refers to the data-taking times after removing low-quality data periods. is the average electron lifetime of each data set. and are high voltage values of the cathode and gate. The values and uncertainties of and are determined by likelihood fits to low energy calibration data [20].
@@
The PandaX-4T utilizes a time projection chamber (TPC), with 3.7 tonnes of LXe in the sensitive volume, to measure the precise position, timing, and energy of each event [21].
@@
Electrons drift upward under the influence of a downward electric field in the sensitive volume, and the delayed electroluminescence signals () are generated in the gaseous xenon region with a much stronger electric field.
Photons from both signals are collected by arrays of Hamamatsu R11410-23 three-inch photomultiplier tubes (PMTs) at the top and bottom of the TPC, respectively.
Combining the photon pattern on the PMT arrays and the time delay between and , the 3D vertex of a given event can be deduced.
@@
The use of is to avoid the influence of potential saturation and dead channel effects from the top PMTs.
Two detector parameters in Eqn. 1 are determined via calibrations [20, 1]: is the photon detection efficiency for , and is the product of electron extraction efficiency and the single-electron gain of .
For this analysis, we utilize the data from Run0 (Nov 28, 2020 - Apr 16, 2021) and Run1 (Nov 16, 2021 — May 15, 2022).
The data is divided into subsets based on different electron drift lifetimes, liquid levels, cathode voltages, and gate voltages.
Detector operating conditions are summarized in Table 1.
The fiducial masses optimized to reject external backgrounds from detector materials are tonne and tonne for Run0 and Run1, respectively.
The signal detection efficiency for Run0 (Run1), including reconstruction efficiency, quality cuts efficiency, and efficiency of ROI, is estimated by both the data-driven method and waveform-simulation method [20, 22].
The ROI is defined with a lower threshold of more than 2 photoelectrons (PE) for and an upper limit of 30 keV in .
Furthermore, a 99.5% ER acceptance cut in the space of log v.s. is applied,
which can reject most of the NR events from neutrons, the coherent elastic scattering of solar 8B neutrino on xenon nucleus, as well as background originating from radon plate-out on the surface of the polytetrafluoroethylene reflector of the TPC.
The overall efficiency in the ROI is estimated to be () for Run0 (Run1), and the energy dependence is shown in Fig. 1.
@@
External detector backgrounds originate from detector components such as stainless steel vessels and PMTs.
Internal backgrounds refer to radioisotopes of xenon and other radio-impurities dissolved in the target.
Table 2: Summary of background expectations and best-fit of the background-only hypothesis of each component for Run0 and Run1. Systematic uncertainties from the signal model and detection efficiencies are included in the uncertainties presented.
@@
The concentration of K, Th, and U of CJPL lab environment are measured at Bq/kg level [24].
With an effective water shielding of at least 4 m in each direction [21], the contribution from surrounding rocks is suppressed to a negligible level.
The concentrations of 232Th and 238U in ultra-pure water shielding itself are measured to be ppt (part per trillion) and are expected to contribute /keV/kg/day background level [25] in the ROI.
Solar neutrinos produced via proton-proton fusion and 7Be electron capture are calculated using standard solar model, three-flavor neutrino oscillation, standard model neutrino-electron scattering, as well as updated treatment of the xenon atomic effects [26, 27, 28] with a 10% uncertainty [29].
The background from gamma rays of detector materials is calculated based on the ER spectrum in the MeV energy range, where characteristic full-absorption peaks can help better constrain the radioactivities of different components in situ [30].
The dominant contributions are from the stainless steel vessels and PMTs.
The external backgrounds in the ROI contribute a flat-shape according to a custom Geant4-based simulation software package BambooMC [31, 32].
The expected contribution to the background in the ROI is 49 3 and 112 10 events for Run0 and Run1, extrapolated from the results of Ref. [30] with the help of BambooMC.
@@
The concentration of 222Rn is determined from the number of decay events and the values for Run0 and Run1 are Bq/kg and Bq/kg respectively.
@@
A depletion effect, i.e., decreased concentration of longer-lived daughter nuclide when compared with the parent in LXe, has been reported [33].
@@
For 220Rn decay chain, the dominating background comes from 212Pb, which is estimated to be Bq/kg with the depletion effect taken into account, using the level of its daughter 212Po -decays as the proxy.
@@
The krypton concentration in Run0 (Run1) is 0.52 0.27 (0.94 0.28) ppt, which is estimated with the coincidence of 85Kr decay to metastable state 85mRb and the subsequent deexcitation to ground state 85Rb.
Assuming an abundance of , the number of 85Kr events within our ROI is 94 for Run0 and 308 for Run1.
The predominant factor contributing to the uncertainty is the limited statistics of - coincidence events, which are 4 events in Run0 and 12 events in Run1.
@@
The half-life of 136Xe’s is measured in Ref. [30], leading to an estimated background contribution of 36.9 2.5 events for Run0 and 66.2 5.9 events for Run1.
Double electron capture events of LL-shell of 124Xe with an expected energy of approximately 10 keV are estimated based on the measured half-life from XENON1T [36].
The expected number of events for Run0 and Run1 is 2.3 0.6 and 4.0 1.1, respectively.
The decay from L-shell electron capture (5.2 keV) of 127Xe also contributes to our ROI, which can be determined by the 408 keV peak (K-shell capture mixed with the 375 keV -ray from 127I).
In addition to the physical background, accidental coincidence (AC) events resulting from the random pairing of isolated- and isolated- signals also contribute to the background.
These spurious events are estimated to be 7.6 2.4 and 7.1 2.3 in Run0 and Run1, respectively [37].
The background levels do not scale with the exposure due to different cut efficiencies from the two runs.
@@
The best fit of each background component after a background-only fit is shown in Fig. 2, with the number of events also summarized in Table 2.
@@
The average fitted ER background rate within the ROI (including efficiency) is and events/tonne/year/keV in Run0 and Run1, respectively.
The data are consistent with the background-only model with a -value of 0.17.
Figure 3: The 90% C.L. upper limits of couplings of different NERS. (a) for ABC axion; (b) for 57Fe axion; (c) for Primakoff axion; (d) for solar neutrino with anomalous magnetic moment; (e) for ALPs; (f) mixing angle for dark photons; (g)(h) cross-sections of light fermionic DM- conversion for the vector and axial-vector interactions. In (g)(h), the limits (blue solid line) and sensitivity median (blue dashed line) of Run0 published results [7], denoted as PandaX-4T (2022), are also overlaid.
@@
NERS hypotheses are tested with the background-plus-signal fits using the PandaX-4T Run0 and Run1 data.
The best fit of the ABC axions hypothesis yields events (ranged to positive-only), while the best fits for axions originating from the Primakoff conversion and 57Fe are null, assuming an axion mass of keV/.
@@
ALPs signals, as well as those from dark photons and light fermionic DM- conversion, exhibit a characteristic peak structure that correlates with their mass.
The mass scanning for ALPs and dark photons is conducted within the range of 1 to 25 keV/, while for light fermionic DM- conversion, the range is from 10 to 180 keV/.
The data exhibit an upward fluctuation in the 3 - 5 keV range for both Run0 and Run1, which results in a local significance of 1.8 for ALPs and dark photons at a mass of 3.3 keV/, and 1.6 for light DM at a mass of 60 keV/ (peak of recoil energy at 4.1 keV) relative to the background-only hypothesis [58].
@@
The upper limits with 90% confidence level (C.L.) of axion couplings (), neutrino magnetic moment , kinematic mixing of dark-to-real photon and light fermionic DM- conversion cross-sections () are shown in Fig. 3.
@@
For conservativeness, limits due to upward fluctuation of data are reported as they are, but those due to downward fluctuation are power-constrained to -1 of the sensitivity band [60].
@@
For axion produced by the 14.4 keV 57Fe transition, are constrained to less than for keV/,
which as competitive as the constraints from XENONnT [8].
The axion-photon coupling is constrained to GeV-1 for keV/, which represents one of the most competitive limits on from a terrestrial experiment, and ruling out [54, 55] and DFSZ [56, 57] axions above a few eV/.
The upper limit on the anomalous magnetic moment of solar neutrino is .
@@
Our sensitivities for light fermionic DM- conversion are notably improved compared to the previous best results in Ref. [7], except at a DM mass around 60 keV/ and between 150 - 180 keV/ due to the upward data fluctuation at around 3 keV and 24 - 25 keV.
@@
In particular, we report leading limits on solar 57Fe axions and light fermionic DM- conversion.
We also report the most stringent limits for the light fermionic DM- conversion for most of the mass range of 10 - 180 keV/.
@@
This project is supported in part by grants from National Science Foundation of China (Nos. 12090060, 12090063, U23B2070), a grant from the Ministry of Science and Technology of China (Nos. 2023YFA1606200, 2023YFA1606201), and by Office of Science and Technology, Shanghai Municipal Government (grant No. 22JC1410100, 21TQ1400218). We thank for the support by the Fundamental Research Funds for the Central Universities. We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Hongwen Foundation in Hong Kong, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.

### v2

@@
zhou_xp@buaa.edu.cnSchool of Physics, Beihang University, Beijing 102206, China
@@ Abstract
We use the large liquid xenon detector PandaX-4T to search for novel electronic recoil signals induced by solar axions, neutrinos with anomalous magnetic moment, axionlike particles, dark photons, and light fermionic dark matter.
A detailed background model is established using the latest datasets with 1.54 exposure.
@@
In addition to the traditional DM-induced nuclear recoil (NR) events, particles from many new physics models can produce novel electronic recoil signals (NERS), which have attracted increasing attention recently [6, 7, 8, 9].
Axions [10, 11, 12] can be produced in the Sun via the ABC effect (atomic recombination and deexcitation, bremsstrahlung, and Compton scattering processes) [13, 14], the Primakoff conversion [15], and the nuclear magnetic transitions [16], related to the axion-electron coupling , axion-photon coupling , and effective axion-nucleon coupling , respectively.
The corresponding effective Lagrangian is given by [17]:
where is the axion field, the electron, the photon, and the nucleon isospin doublet .
and are model-dependent isoscalar and isovector axion-nucleon couplings and is the Pauli matrix.
The Sun also produces a vast number of neutrinos, which may carry anomalous magnetic moment due to the Majorana nature of neutrino or exotic new physics beyond the standard model (SM) [18].
With their kinetic energy governed by the Sun’s temperature, these particles can produce electronic recoil (ER) events with typical energy from keV to tens of keV.
Within this energy range, alternative light dark matter models such as axionlike particles (ALPs), dark photons [19], or light fermionic DM converting into active neutrinos (DM-) [20, 21] can also manifest themselves with characteristic ER spectra.
ALPs interact with SM particles similarly to axions, while dark photons interact through kinetic mixing with SM photons.
The vector and axial-vector interactions among the dimension-six effective operators of DM- conversion can be probed experimentally. Considering the atomic effect, the differential cross section with respect to the ionized electron energy is
where and are the principal quantum number and angular momentum.
— is the energy of outgoing neutrino and is the energy of the accompanying x-ray, which has the same energy as the binding energy of state .
is the particle scattering amplitude and
is the atomic factor [22].
The masses of electron and DM particle are denoted as and , respectively.
This Letter reports an updated NERS search using the low-energy ER data below 30 keV from PandaX-4T, an LXe detector operating in the China Jinping Underground Laboratory (CJPL), with a total exposure of 1.54 tonyr.
@@
We leave detailed production and detection cross sections and expected NERS spectra in the Supplemental Material [23] for brevity.
To set the stage, the signals of ABC solar axions and the DM- conversion of keV/ are shown in Fig. 1, with spectra at the low and high end of the region of interest (ROI).
The keys to this analysis lie in the understanding of the detector responses, as well as robust prior knowledge of the background from data outside the ROI.
Figure 1: Theoretical (shaded) and expected energy spectra (solid for Run0, dashed for Run1) for ABC axion with a coupling =3.1 (red shaded for theoretical, yellow unshaded for expected) and DM- conversion with = 170 keV and cm2 (blue shaded for theoretical, purple unshaded for expected).
@@
Table 1: Operating conditions of Run0 and Run1. Duration represents the calendar days, and live time refers to the data-taking time after removing low-quality data periods. is the average electron lifetime of each dataset. and are voltages of the cathode and gate. and values are determined by likelihood fits to low energy calibration data [41].
@@
The PandaX-4T utilizes a time projection chamber (TPC), with 3.7 ton of LXe in the sensitive volume, to measure the precise position, timing, and energy of each event [42].
@@
Electrons drift upward under the influence of a downward electric field in the sensitive volume, and the delayed electroluminescence signals () are generated in the gaseous xenon region.
Photons from both signals are collected by arrays of Hamamatsu R11410-23 three-inch photomultiplier tubes (PMTs) at the top and bottom of the TPC.
Combining the photon pattern on the PMT arrays and the time delay between and , the three-dimensional vertex of a given event can be deduced.
@@
Using avoids the influence of potential saturation and dead channel effects from the top PMTs.
is the photon detection efficiency for .
is the product of electron extraction efficiency and the single-electron gain of .
Both parameters are determined via calibrations [41, 1].
For this analysis, we utilize the data from Run0 (Nov. 28, 2020 - Apr. 16, 2021) and Run1 (Nov. 16, 2021 — May 15, 2022), which are divided into subsets based on different electron lifetimes, liquid levels, cathode voltages, and gate voltages, as summarized in Table 1.
The fiducial masses optimized to reject external backgrounds from detector materials are and ton for Run0 and Run1, respectively.
The signal detection efficiency for Run0 (Run1), including reconstruction efficiency, quality cuts efficiency, and ROI efficiency, is estimated by both the data-driven method and waveform-simulation method [41, 43].
The low threshold of the ROI is 2 photoelectrons (PE) for , and the upper limit is 30 keV in .
A 99.5% ER acceptance cut in the space of log vs is applied, which can reject more than 66% (99%) of the NR events from neutrons (the coherent elastic scattering of solar 8B neutrino on xenon nucleus) and 95% of NR-like background originating from radon plate-out from the polytetrafluoroethylene reflector surface in the TPC.
The overall efficiency is estimated to be () for Run0 (Run1), and the energy dependence is shown in Fig. 1.
Above 7 keV, the efficiency plateaus and the average value is () for Run0 (Run1).
@@
External backgrounds originate from detector components such as PMTs.
Internal backgrounds refer to radioisotopes of xenon and other radio impurities dissolved in the target.
Table 2: Summary of background expectations and best fit of the background-only hypothesis of each component for Run0 and Run1. Systematic uncertainties from the signal model and detection efficiencies are included in the uncertainties presented.
@@
Solar neutrinos produced via proton-proton fusion and 7Be electron capture are calculated using standard solar model, three-flavor neutrino oscillation, standard model neutrino-electron scattering, as well as updated treatment of the xenon atomic effects [45, 46, 47].
A 10% uncertainty is used based on Borexino measurements [48].
The concentration of K, Th, and U of the CJPL lab environment are measured at Bq/kg level [49].
With at least 4 m of water shielding in each direction [42], the contribution from surrounding rocks is suppressed to a negligible level.
The concentrations of 232Th and 238U in ultra-pure water shielding itself are measured to be ppt (part per trillion) and are expected to contribute events/ton/yr/keV [50] in the ROI.
The background from rays of detector materials is calculated based on the ER spectrum in the MeV energy range, where characteristic full-absorption peaks can help constrain the radioactivities of different components in situ [51].
The external backgrounds contribute a flat shape in the ROI according to a custom geant4-based simulation package BambooMC [52, 53], with the dominant contributions from the stainless steel vessels and PMTs.
The expected background counts in the ROI are and events for Run0 and Run1, extrapolated from the results of Ref. [51] with the help of BambooMC.
@@
The concentration of 222Rn is () Bq/kg for Run0 (Run1), determined from the number of decay events.
@@
A depletion effect, i.e., decreased concentration of longer-lived daughter nuclide compared to the parent in LXe, has been reported [54].
@@
For 220Rn decay chain, the dominating background comes from 212Pb, which is estimated to be Bq/kg with the depletion effect taken into account, using the level of its daughter 212Po decays as the proxy.
@@
The krypton concentration in Run0 (Run1) is 0.52 0.27 (0.94 0.28) ppt, estimated with the coincidence of 85Kr decay to metastable state 85mRb and the subsequent deexcitation to ground state 85Rb.
The predominant contribution to the uncertainty is the limited statistics of - coincidence, which are 4 events in Run0 and 12 events in Run1.
Assuming an isotopic abundance of , 94.2 (308.1) 85Kr events are in our ROI for Run0 (Run1).
@@
The half-life of 136Xe’s double beta decay is measured in Ref. [51], leading to an estimated background contribution of events for Run0 and events for Run1.
Double electron capture events of LL-shell of 124Xe with an expected energy of approximately 10 keV are estimated based on the theoretical branching ratio and the measured 124Xe half-life from XENON1T [57].
The expected number of events for Run0 and Run1 is and , respectively.
The decay from -shell electron capture (5.2 keV) of 127Xe also contributes to our ROI, which can be determined by the 408 keV peak (-shell capture mixed with the 375 keV -ray from 127I).
In addition to the physical background, accidental coincidence (AC) events resulting from the random pairing of isolated and signals also contribute to the background.
These spurious events are estimated to be and in Run0 and Run1, respectively [58].
The background levels do not scale with the exposure due to different cut efficiencies in the two runs.
@@
The best fit of a background-only model is shown in Fig. 2, with the number of fitted background events summarized in Table 2.
@@
The ROI’s average fitted ER background rate (including efficiency) is and events/ton/yr/keV in Run0 and Run1, respectively.
The data are consistent with the background-only model with a value of 0.16.
Figure 3: The 90% CL upper limits of couplings of different NERS. (a) for ABC axion; (b) for 57Fe axion; (c) for Primakoff axion; (d) for solar neutrino with anomalous magnetic moment ( is the Bohr magneton); (e) for ALPs; (f) mixing angle for dark photons; (g)(h) cross sections of light fermionic DM- conversion for the vector and axial-vector interactions. In (g)(h), the limits (blue solid line) and sensitivity median (blue dashed line) of Run0 published results [7], denoted as PandaX-4T (2022), are also overlaid.
In scenarios (e) to (h), we assume that each type of dark matter candidate constitutes 100% of the local dark matter.
@@
NERS hypotheses are tested with the background-plus-signal fits using the Run0 and Run1 data.
The best fit of the ABC axions hypothesis yields events (ranged to positive only), while the best fits for axions originating from the Primakoff conversion and 57Fe are null, assuming an axion mass of keV/.
@@
ALP signals, as well as those from dark photons and DM- conversion, exhibit a characteristic peak structure that correlates with their mass.
The mass scanning for ALPs and dark photons is conducted within the range of 1 to 25 keV/, while for DM- conversion, the range is from 10 to 180 keV/.
The data exhibit an upward fluctuation in the 3 - 5 keV range for both Run0 and Run1, which results in a local significance of 1.9 for ALPs and dark photons at a mass of 3.3 keV/, and 1.6 for light DM at a mass of 60 keV/ (peak of recoil energy at 4.1 keV) relative to the background-only hypothesis [78].
@@
The upper limits with 90% confidence level (CL) of axion couplings (), , kinetic mixing of dark-to-SM photon and DM- conversion cross sections () are shown in Fig. 3.
@@
For conservativeness, limits due to upward fluctuation of data are reported as they are, but those due to downward fluctuation are power constrained to -1 boundary of the sensitivity band [80].
@@
For axions produced by the 14.4 keV 57Fe transition, are constrained to less than for keV/, which is as competitive as the constraints from XENONnT [8].
The axion-photon coupling is constrained to GeV-1 for keV/, which represents one of the most competitive limits on from a terrestrial experiment, and disfavoring KSVZ [74, 75] and DFSZ [76, 77] axions above a few eV/.
The upper limit on the anomalous neutrino magnetic moment is .
@@
Our sensitivities for DM- conversion are notably improved compared to the previous best results in Ref. [7], except at around 60 keV/ and between 150 - 180 keV/ due to the upward data fluctuation at around 3 keV and 24 - 25 keV.
@@
In particular, we report leading limits on solar 57Fe axions and light fermionic DM- conversion, the latter of which is the most stringent result for most of the mass range of 10 - 180 keV/.
@@
This project is supported in part by grants from National Key R&D Program of China (No. 2023YFA1606200, No. 2023YFA1606201, National Science Foundation of China (No. 12090060, No. 12090061, No. 12105008, No. U23B2070), and by Office of Science and Technology, Shanghai Municipal Government (Grant No. 21TQ1400218, No. 22JC1410100, No. 23JC1410200, No. ZJ2023-ZD-003). We thank the support by the Fundamental Research Funds for the Central Universities, Hongwen Foundation in Hong Kong, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.